# QMol-grid: A MATLAB package for quantum-mechanical simulations in atomic and molecular systems

### v2

@@ 4 Impact
The QMol-grid package offers a versatile suite of quantum simulation techniques for reduced-dimension atomic and molecular models. Its native MATLAB structure facilitates on-the-fly calculations and analyses in time-dependent simulations as well as post-processing, which all can be done using high-level functionalities of MATLAB. Simulation data are organized within handle classes with common interface methods to simplify end-user interaction with the various components of the package. QMol-grid comes with a full documentation, including many script samples that illustrate how one can use the various features. It also includes a series of tutorials to guide new users with setting up calculations, input parameters, and output variables.
@@ 5 Conclusions
The QMol-grid package provides a versatile suite of quantum-mechanical methods at the Schrödinger, Hartree-Fock, and density-functional theory levels of theory for ground- and excited-state calculations, as well as TDSE and TDDFT propagators. Time-propagation schemes provide streamlined access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT). The wave functions and Kohn-Sham orbitals are packaged into classes that enable abstract manipulations in the objects, e.g., for ground-state, time propagation, and common observables’ calculations. The object-oriented structure provides a uniform user interface, where input parameters are specified as pairs of parameter-name/parameter-value (in arbitrary order and case insensitive). Output results are stored in the objects and can be recovered using standard object-oriented dot notation – see the tutorials for examples.
@@ Acknowledgements
FM thanks M.B. Gaarde, K. Lopata, and K.J. Schafer for enlightening discussions, suggestions, and support throughout the development of the package. The original development of the QMol-grid package, and its (TD)DFT features, was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award No. DE-SC0012462. Addition of the (TD)SE features was supported by the National Science Foundation under Grant No. PHY-2207656

### v3

@@ 4 Impact
    The |QMol-grid|package offers a versatile suite of quantum simulation techniques for reduced-dimension atomic and molecular models. Its native MATLAB structure facilitates on-the-fly calculations and analyses in time-dependent simulations as well as post-processing, which all can be done using high-level functionalities of MATLAB. Simulation data are organized within handle classes with common interface methods to simplify end-user interaction with the various components of the package. |QMol-grid|comes with a full documentation, including many script samples that illustrate how one can use the various features. It also includes a series of tutorials to guide new users with setting up calculations, input parameters, and output variables.
@@ 5 Conclusions
    The |QMol-grid|package provides a versatile suite of quantum-mechanical methods at the Schrödinger, Hartree-Fock, and density-functional theory levels of theory for ground- and excited-state calculations, as well as TDSE and TDDFT propagators. Time-propagation schemes provide streamlined access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT). The wave functions and Kohn-Sham orbitals are packaged into classes that enable abstract manipulations in the objects, e.g., for ground-state, time propagation, and common observables’ calculations. The object-oriented structure provides a uniform user interface, where input parameters are specified as pairs of parameter-name/parameter-value (in arbitrary order and case insensitive). Output results are stored in the objects and can be recovered using standard object-oriented dot notation – see the tutorials for examples.
@@ Acknowledgements
    FM thanks M.B. Gaarde, K. Lopata, and K.J. Schafer for enlightening discussions, suggestions, and support throughout the development of the package. The original development of the |QMol-grid|package, and its (TD)DFT features, was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award No. DE-SC0012462. Addition of the (TD)SE features was supported by the National Science Foundation under Grant No. PHY-2207656